### v2

# Networks with many structural scales: a Renormalization Group perspective

Abstract
Scale invariance profoundly influences the dynamics and structure of complex systems, spanning from critical phenomena to network architecture. Here, we propose a precise definition of scale-invariant networks by leveraging the concept of a constant entropy loss rate across scales in a renormalization-group coarse-graining setting. This framework enables us to differentiate between scale-free and scale-invariant networks, revealing distinct characteristics within each class. Furthermore, we offer a comprehensive inventory of genuinely scale-invariant networks, both natural and artificially constructed, demonstrating, e.g., that the human connectome exhibits notable features of scale invariance. Our findings open new avenues for exploring the scale-invariant structural properties crucial in biological and socio-technological systems.
The network paradigm effectively captures essential attributes of real-world complex systems, offering a natural framework for studying entangled interconnected systems across disciplines like neuroscience Fornito et al. (2016), ecology Bascompte (2007), and epidemiology Pastor-Satorras and Vespignani (2001), among others Newman (2003). Understanding the evolutionary dynamics of complex networks, as they adapt their connectivity patterns to achieve diverse goals, is crucial to understanding their long-term stability or other features influencing functional roles and performance Fortuna et al. (2011); Villegas et al. (2020). Notably, amidst the multitude of potential network structures, one organization ubiquitously arises in natural systems: the scale-free architecture Barabási and Albert (1999); Newman (2010); Barabási (2016); Caldarelli et al. (2002).
Scale-free networks manifest a distribution of node connectivities $k$ that decays as a power-law $P(k)\propto k^{-\gamma}$ for large values of $k$ Barabási and Albert (1999); Broido and Clauset (2019). In statistical physics, power-law behavior is the hallmark of scale invariance and scaling behavior, implying no significant characteristic value of the analyzed quantity Newman (2005); Sornette (2006); Muñoz (2018). For scale-free networks, there is no typical scale in the degree of connectivity apart from natural cut-offs. However, the network community has been intrigued by the possibility of discerning, through careful statistical analyses Clauset et al. (2009); Broido and Clauset (2019), including finite-size effects Serafino et al. (2021), whether empirical networks genuinely exhibit bona-fide scale invariance or only appear to. From a theoretical standpoint, addressing this question calls for designing a renormalization group (RG) approach Wilson (1979); Amit and Martin-Mayor (2005); Caldarelli et al. (2024); Kardar (2007). In this framework, a fixed point of the RG transformation denotes a state or system whose characteristics remain unchanged under appropriate scale transformations achieved through iterated coarse-graining. RG fixed points are inherently linked to universal scaling laws governing the system, so systems at the same fixed point share identical scaling features and belong to the same universality class Amit and Martin-Mayor (2005); Kardar (2007). Hence, a key implication of RG theory is the categorization of numerous seemingly disparate physical systems and dynamical models into a relatively compact set of universality classes at criticality, either in or out of thermal equilibrium Hohenberg and Halperin (1977); Ódor (2004). Elucidating the relevant ingredients that give rise to a particular universality class offers valuable insights into the fundamental mechanisms underpinning complex systems and their key features.
Given the lack of a natural Euclidean embedding for complex networks, traditional length-scale transformations using translational invariance were deemed unfeasible until recently Radicchi et al. (2008); Rozenfeld et al. (2007). In particular, small-world effects, characterized by short path lengths between nodes, further complicate block identification or affine transformations Cohen and Havlin (2003); Watts and Strogatz (1998); Rozenfeld et al. (2007, 2010). Nevertheless, various techniques have been proposed to tile the network, exploiting specific scale-invariant properties as a function of the number of links along any shortest path between two nodes Radicchi et al. (2008). These include box-covering techniques Song et al. (2005); Rozenfeld et al. (2010), spectral partitioning Gfeller and De Los Rios (2007), and hyperbolic geometry embeddings, which offer a promising novel approach for understanding particular complex network structures and dynamics Serrano et al. (2008); García-Pérez et al. (2018). However, fully characterizing scale-invariant properties in real networks remains an open challenge.
A novel approach, the Laplacian Renormalization Group (LRG), generalizes statistical mechanics RG to graphs, providing a comprehensive framework for coarse-graining heterogeneous systems in both real and momentum space Villegas et al. (2023a). This approach uses diffusive dynamics on networks to gradually remove the smallest fine-grained structural scales, eliminating the contribution from large eigenvalues. Specifically, the LRG allows for the redefinition of effective time-dependent Laplacian and adjacency matrices at coarser scales. Due to its general validity, it comes natural to explore structural scale invariance within the LRG framework, as evidenced by the observation of a time-independent rate of entropy loss Villegas et al. (2023a); Klemm (2023). This leads to the fundamental question of whether non-Euclidean architectures, like general networks, can be intrinsically scale-invariant. Two key questions arise. Can we identify benchmark classes of network architectures exhibiting genuine scale-invariant behavior and, thus, define an analog to universality classes? Can we detect and quantify this kind of self-similarity in real-world networks? Here, we provide a clear definition and compilation of scale-invariant networks, emphasizing that scale-freeness, i.e., a power law degree distribution, does not imply structural self-similarity, and vice versa. Finally, we analyze actual brain networks and highlight their scale-invariant features.
Laplacian Renormalization Group (LRG).
The introduction of a multi-scale statistical mechanical approach, the LRG, allows the detection of different scales within a network Villegas et al. (2023a, b). The LRG is based on the time-evolution operator $e^{-\tau\hat{L}}$ of the diffusion or heat equation, where $\hat{L}=\hat{D}-\hat{A}$ is the Laplacian operator, $\hat{A}$ is the adjacency matrix, and $\hat{D}$ is the degree matrix Villegas et al. (2022). Using this operator and denoting the Laplacian eigenvalues as $\lambda_{i}$ with $i=1,...,N$ (all real and positive Chung (1997); Donetti et al. (2006)), one can define the Laplacian density matrix De Domenico and Biamonte (2016),
$\hat{\rho}(\tau)=\frac{e^{-\tau\hat{L}}}{Z\equiv Tr(e^{-\tau\hat{L}})}=\frac{e^{-\tau\hat{L}}}{\sum_{i=1}^{N}e^{-\lambda_{i}\tau}}\,,$ (1)
and develop a "canonical" description of heterogeneous networks in analogy with statistical mechanics Villegas et al. (2023a, 2022). Here, $\hat{L}$ formally acts as a Hermitian Hamiltonian, $Z(\tau)$ is the partition function, and $\tau$ serves as a control scale parameter akin to the inverse temperature. In this way, as the resolution scale $\tau$ is increased, the contribution of large eigenvalues to $\hat{\rho(\tau)}$revealing fine structure- is progressively removed, allowing for an effective network coarse-graining  [^1] This idea is also at the basis of different existing community detection algorithms Villegas et al. (2023b); Arenas et al. (2006); Donetti and Muñoz (2004); Donetti et al. (2005). Within this formalism, one can compute the network entropy as $S(\tau)=-\mbox{Tr}[\hat{\rho}(\tau)\log\hat{\rho}(\tau)]$, so that $S(\tau)=\tau\langle\lambda\rangle_{\tau}+\ln Z(\tau)$, where $\langle\lambda\rangle_{\tau}\equiv\langle\hat{L}\rangle_{\tau}=\nicefrac{{\stackrel{{\scriptstyle[}}{{i}}=1]{N}{\sum}\lambda_{i}e^{-\tau\lambda_{i}}}}{{Z(\tau)}}$. In particular, $S$ runs from $S=\ln N$ at $\tau=0$, the segregated (high temperature) regime to $S=0$ at $\tau\rightarrow\infty$, the integrated (low temperature) regime. More specifically, one can define the network entropic susceptibility or "heat capacity" Villegas et al. (2022)
$C(\tau)\equiv-\frac{dS}{d\log\tau},$ (2)
which describes the rate of entropy loss —i.e., the rate at which the complexity of the network shrinks upon coarse-graining— or the rate of information acquired about the network structure during diffusion dynamics at scale $\tau$. Recalling statistical physics, peaks of $C$ (diverging in the infinite size limit) are associated with structural phase transitions. Thus, we can analyze $C$ at varying $\tau$ to investigate the network multi-scale organization, detecting scales where entropy changes more significantly due to structural transitions. This description permitted, e.g., the detection of the network information core and its related structural/diffusive transitions Villegas et al. (2022) and led to naturally extending the RG to heterogeneous networks Villegas et al. (2023a). Here, it allows us to define informationally scale-invariant networks as graphs whose entropy-loss rate $C(\tau)$ takes a constant value $C_{0}>0$ across scales or, at least, within a sufficiently broad diffusion-time interval, thus being (exactly or approximately) scale invariant.
We demonstrate that this definition of informational scale-invariance holds if and only if the Laplacian spectral density follows $w(\lambda)\sim\lambda^{\gamma}$, considering the spectral density as a continuum distribution in the infinite network-size limit. Using Eq.(2) and the definition of $S(\tau)$ one derives $C(\tau)=-\tau^{2}\frac{d\langle\lambda\rangle_{\tau}}{d\tau}$, and setting $C(\tau)=C_{0}$ as a constant, one gets $\langle\lambda\rangle_{\tau}=C_{0}/\tau$. The only solution to this equation is a Laplacian spectral density $w(\lambda)\sim\lambda^{\gamma}$, which allows expressing $C_{0}$ as a function of the exponent $\gamma$, specifically $C_{0}=\gamma+1$. This implies that macroscopic properties —stemming from vanishing Laplacian eigenvalues— of a network with a power-law spectrum remain invariant under LRG scale transformations. Thus, network scale-invariance shifts from a power-law degree distribution to a power-law in the spectral density. Note that the exponent of the Laplacian spectral density, $\gamma$, can be related to the spectral dimension, $d_{s}$, which is a global property of the graph, related, e.g., to the infrared singularity of the Gaussian process Burioni and Cassi (1996) and that has been shown to provide a robust generalization of the standard dimension for networks Cassi (1992); Burioni and Cassi (1996). As $d_{s}/2=\gamma+1$ Burioni and Cassi (1996), one concludes that $C_{0}$ is constant and equals half of the graph’s spectral dimension: $C_{0}=d_{s}/2$. Thus, measuring the plateau value of the heat capacity of a scale-invariant network effectively determines its dimensionality.
For finite-sized networks, scale invariance may only be approximate and confined to a finite-scale interval. This means that $C(\tau)$ cannot remain constant across all scales: it decays from its plateau value $C_{0}$ to 00 for sufficiently large values of $\tau$. To estimate this cut-off, we have to consider that the smallest non-zero eigenvalue $\lambda_{2}$ of $\hat{L}$, also called "spectral gap" or "Fiedler eigenvalue" $\lambda_{F}$, Fiedler (1989); Chung et al. (2003), is the one providing the slowest decaying contribution in time to Eq.(1), therefore determining the asymptotic time decay of $C(\tau)$. Starting from Eq.(2) and imposing that the Laplacian spectrum $\omega(\lambda)$ integrates to $N$, we get $\lambda_{F}\sim N^{\nicefrac{{-2}}{{d_{s}}}}$ for large $N$ Das (2004); Dorogovtsev and Mendes (2022). From this, it follows (see Supplemental Material, SM SM ),
${\small C\left(\frac{\alpha}{\lambda_{F}}\right)\simeq\frac{\Gamma\left(\frac{d_{s}}{2}\right)\Gamma\left(\frac{d_{s}}{2}+2,\alpha\right)-\Gamma^{2}\left(\frac{d_{s}}{2}+1,\alpha\right)}{\Gamma^{2}\left(\frac{d_{s}}{2}\right)},}$ (3)
[…]
Laplacian Random-Walk RG.
The LRG method relies on a “heat-like" diffusion process defined by $\hat{L}$. However, we wonder whether the results obtained for scale-invariance and graph dimensionality remain valid when using the Laplacian random-walk operator, $\hat{L}_{RW}=\hat{D}^{-1}\hat{L}$ instead of $\hat{L}$. $\hat{L}_{RW}$ describes the time-discrete dynamics of a RW moving from a node to a neighbor with uniform probability and represents the transition matrix for RW dynamics on a graph Burioni and Cassi (2005); Chung (1997). Although $\hat{L}_{RW}$ is not symmetric, it is equivalent to the symmetric operator $\hat{L}_{sym}=D^{1/2}\hat{L}_{RW}\hat{D}^{-1/2}$ Chung (1997) with all its eigenvalues real and satisfying $0\leq\mu_{i}\leq 2$. We can, thus, reformulate the LRG for the RW dynamics by substituting $\hat{L}$ with $\hat{L}_{sym}$, renaming the related heat capacities as $C_{L}$ and $C_{RW}$, respectively. In fact, $\hat{L}_{RW}$ provided the original definition of the RW spectral dimension, $d_{s}^{RW}$, related to the first-return time distribution of the random-walker $P_{0}(t)$ Burioni and Cassi (1996); Cassi and Fabbian (1999) by the scaling $P_{0}(t)\propto t^{-d_{s}^{RW}/2}$. This extends the behavior on $d-$dimensional lattices where $P_{0}(t)\propto t^{-d/2}$ Alexander and Orbach (1982). While $\hat{L}$ or $\hat{L}_{RW}$ strictly coincide in regular networks like lattices, both operators unravel heterogeneous network structures in different ways Villegas et al. (2023b) and, to our knowledge, there is no proof that $d_{s}=d_{s}^{RW}$ for generic heterogeneous networks.
LRG non-trivial fixed points.
Subsequently, we tackle the following lingering questions: What are the current non-trivial fixed points of the LRG, i.e., families or "universality classes" of scale-invariant networks? Are scale-invariant networks necessarily scale-free? Do results depend on the choice of the operator?
Let us remark that regular lattices, with a degree distribution $P(\kappa)=\delta(\kappa-\kappa_{0})$, represent the simplest case of scale-invariant structures lacking scale-free properties (see SM SM ). However, our main focus is categorizing heterogeneous and stochastic scale-invariant networks.
Figure 1: Trees. Heat capacity versus diffusion time for: (a)-(b) Random trees and (c)-(d) Barabasi-Albert networks with $m=1$. We have used $\hat{L}$ in (a) and (c) and $\hat{L}_{RW}$ in (b) and (d). Insets show the scaling of $\tau_{F}$ as a function of the network size $N$. Orange dashed lines and solid black lines represent the theoretical expectation, $d_{s}^{RT}=\nicefrac{{4}}{{3}}$ and $d_{s}^{BA}=2$. All curves have been averaged over $10^{3}$ independent realizations. Different colors represent different sizes $N=\{1,2.5,4,8,16\}\times 10^{3}$.
The first category of self-similar networks comprises trees, i.e., connected loopless networks. It is known that the spectral dimension of random trees, with minimal branching ratio $b_{min}=1$, depends upon the first two moments of the degree distribution $P(\kappa)$ Donetti and Destri (2004); Burda et al. (2001). Specifically, $d_{s}=4/3$ if $\langle\kappa^{2}\rangle$ is finite, while the problem remains open when $\langle\kappa^{2}\rangle$ diverges Donetti and Destri (2004); Destri and Donetti (2002); Dorogovtsev et al. (2003). Here, we examine specific cases within these two classes, with the lenses of LRG: ordinary random trees (RT) and Barabási-Albert (BA) networks where new nodes attach preferentially to existing ones, forming $m=1$ edges Barabási and Albert (1999). As reported in Fig.1(a), $C_{L}$ shows a plateau corresponding to the theoretically known spectral dimension for random trees, $d_{s}^{RT}=\nicefrac{{4}}{{3}}$ Destri and Donetti (2002) and follows the finite-size scaling condition, Eq.(3). Also, we observe that $\hat{L}_{RW}$ does not alter this value (see Fig.1(b)), despite specific differences in the shape of $C$, as the local peak at short times in Fig.1(b).
Instead, BA networks with $m=1$ are trees with diverging $\langle\kappa^{2}\rangle$. As shown in Fig.1(c), they have a constant heat capacity $C_{L}$ corresponding to a spectral dimension $d_{s}=2$. Although $C_{RW}$ grows in the intermediate regime, it also becomes flat for asymptotic times and large network sizes (see Fig.1), confirming scale-invariance with $d_{s}^{RW}=d_{s}=2$. This same spectral dimension is also obtained for Bethe lattices with coordination number $z\geq 3$ (see Erzan and Tuncer (2020) and SM SM ). Furthermore, BA networks with $m>1$ show no sign of scale invariance (see SM SM ), proving that the archetypes of scale-free networks are not scale-invariant. This prevents preferential attachment from generating self-similar networks with $d_{s}>2$.
Second, we consider networks that, contrary to trees, have a non-vanishing clustering coefficient. In particular, the Dorogovstev-Golstev-Mendes Dorogovtsev et al. (2002) networks and their generalization: (u,v)-flowers Dorogovtsev et al. (2002); Rozenfeld et al. (2007, 2009). These deterministic fractal-like structures grow by iteratively replacing the link between two nodes with two paths of lengths $u$ and $v$ (see SM SM ). Diameter-based analyses, such as calculating the Hausdorff dimension of these networks lead to an apparent paradox: all $(1,v)$ flowers are infinite-dimensional and exhibit anomalous scaling functions Rozenfeld et al. (2007); Dorogovtsev et al. (2002); Peng and Agliari (2017). This paradox stems from the lack of network embedding in an Euclidean space. Conversely, as shown in Figs.2(a), (b) and SM SM , all flowers exhibit well-defined spectral dimensions Rozenfeld et al. (2007), $d_{s}=\log(\emph{u}+\emph{v})/\log(\emph{u}\emph{v})$ changing from $d_{s}\approx 3.17$ to $d_{s}=1$.
Figure 2: Clustered networks. RW heat capacity versus diffusion time for: (a) $(1,2)$ flowers with $s=6,8,9,10$ and $11$, hierarchical levels (b) $(2,2)$ flowers with $s=5,6,7$, and $8$. For $(u,v)$ flowers, the number of nodes is $N_{s}=\left(\frac{w}{w-1}+\frac{w-2}{w-1}w^{s}\right)$, with $w=u+v$. (c) KH with $m=2$ and (d) KH with $m=3$ of sizes $N=\{1,2.5,4,8\}\times 10^{3}$. Insets show the scaling $\tau_{F}$ as a function of the system size. Orange dashed lines and solid black lines represent the theoretical expectation. All curves have been averaged over $10^{3}$ independent realizations.
Fig.2(b) sheds light on their differences: $(u,v)$ flowers with $u>1$ exhibit characteristic oscillations emerging from the discrete nature of the recursive network growth process (much as log-periodic oscillations in standard fractal analyses of deterministic structures, as stem from their discrete-scale invariance Sornette (2006); Vicsek et al. (1994)). Thus, we demonstrate that the method detects the presence of discrete scale invariance on networks.
Third, to explore whether self-similarity can exist in stochastic scale-free graphs with small-world properties Rozenfeld et al. (2007), we consider the Kim and Holme (KH) model Holme and Kim (2002). This requires modifying BA networks by introducing a probability $p$ for each newly added node to form a triangle with existing nodes (see SM SM ). The resulting networks show a power-law degree distribution with high clustering coefficient. LRG analysis of KH networks reveals scale-invariance only when $p=1$. As shown in Figs.2(c) and (d), $C_{RW}$ (and $C_{L}$ SM ) shows a plateau with spectral dimension from $d_{s}=2.57(1)$ to $d_{s}=3.65(1)$ as $m$ increases (see SM SM ).
Fourth, we have investigated the behavior of networks with a built-in hierarchical structure. In particular, we have analyzed the Dyson graph Dyson (1969); Agliari and Tavani (2017): a fully-connected deterministic graph with hierarchically organized link weights, resulting in a tunable spectral dimension $d_{s}=2/(2\sigma-1)$ where $\sigma$ is a scaling parameter controlling the weight strength at every scale (see SM SM ). Fig.3(a) shows the constant plateau of $C_{L}$ for a Dyson network with $\sigma=0.85$, consistent with the theoretical value. We have also examined hierarchical modular networks (HMNs) that were proposed using inspiration from actual brain networks Moretti and Muñoz (2013). In HMNs, nodes grouped in basal fully-connected modules are recursively coupled with nodes in other moduli, establishing inter-modular links randomly in a hierarchical fractal-like manner (see SM SM ). Fig.3(b) shows their heat capacity: HMNs are scale-invariant for any set of parameters, with spectral dimension in the range $d_{s}\in(1.25,2)$ (see SM SM ).
Figure 3: Hierarchical networks. Heat capacity versus diffusion time for: (a) Dyson graphs with $\sigma=0.85$ and $s=11,13,14,15$ hierarchical levels. The network size is $N=2^{s}$ in both cases. The upper left inset shows the weighted adjacency matrix. (b) HMNs with $m_{0}=3$ and $\alpha=2$ (see SM SM ) for $s=9,10,11,12$ hierarchical levels ($N=2^{s}m_{0}$). Insets show the scaling of $\tau_{F}$ as a function of $N$. Orange dashed and solid black lines represent the theoretical expectation. (c) The HC for different thresholds (see legend). (d) The scaling of $\tau_{F}$ for coarse-grained versions of the HC with $T=3\cdot 10^{-2}$ and different fixed values of $\tau=10^{-1}$, $10^{0}$, and $10^{1}$. Black dashed line corresponds to $d_{s}=1.9$.
Finally, we have analyzed the Human Connectome (HC) structural brain network Fornito et al. (2016) (see SM SM ). The analysis of $C_{L}$ for the largest available (thresholded) HC network is shown in Fig.3(c), while Fig.3(d) shows the finite-size scaling of $\lambda_{F}$ for reduced versions of the HC. Both analyses support an underlying scale-invariant topology with spectral dimension $d_{s}\sim 1.9(1)$.
Outlook.
The observation of scale-invariance across length and time scales led to the development of RG ideas Kardar (2007); Amit and Martin-Mayor (2005) whose initial implementation took advantage of the geometrical homogeneity and translational invariance to perform scaling of statistical physical systems, computing critical exponents Wilson and Fisher (1972), and leading to a categorization of apparently diverse phase transitions into a few universality classes Hohenberg and Halperin (1977); Ódor (2004).
Extending universality to networks with heterogeneous and non-local connectivity properties has long challenged physicists. Despite remarkable advances have been recently made Song et al. (2005); Serrano et al. (2008); Rozenfeld et al. (2010); García-Pérez et al. (2018); Zheng et al. (2020), fully understanding scale-invariance in generic networks remains unresolved. Initially, network scale-invariance was associated with power-law degree distributions Serafino et al. (2021). However, as we explicitly show, scale-freeness of node degrees is neither necessary nor sufficient for true self-similarity Broido and Clauset (2019). Instead, the LRG Villegas et al. (2023a, 2022), provides a general framework to characterize and classify non-trivial structural LRG fixed points, describing their scaling properties and offering a classification in terms of universality classes. A constant heat-capacity indicates the presence of scale-invariance or self-similarity in graphs, characterized by a constant entropy-loss rate during network coarse-graining. We have linked the constant entropy-loss rate to the network spectral dimension, $d_{s}$ —computed using either $\hat{L}$ or $\hat{L}_{RW}$— confirming its role as a natural generalization of the Euclidean dimension Cassi (1992); Cassi and Fabbian (1999); Burioni and Cassi (1996).
Our LRG analysis identifies regular lattices, trees, clustered networks, and hierarchical networks as fundamental classes of scale-invariant networks. Beyond regular lattices, we highlight (u-v)-flowers and KH networks as unique structures combining local clustering with hubs at all scales to generate self-similar topologies. Trees are bound to $d_{s}\leq 2$, and simply adding hubs appears insufficient to achieve higher-dimension values. Generally, as the dimension increases, it becomes more challenging to generate self-similar heterogeneous networks Rozenfeld et al. (2007). Our findings are promising for studying biological and socio-technological networks; notably, hierarchical networks, including the HC Sporns et al. (2004); Hagmann et al. (2008) which exhibit robust scale invariance with $d_{s}\approx 1.9(1)$, corroborating previous results Zheng et al. (2020). Our findings provide a solid ground for future dynamical RG theoretical calculations on these structures and for identifying new signatures of scale invariance in real-world networks.
Acknowledgements.
Acknowledgments
The authors thank T.Gili, D. Cassi, and R. Burioni for very useful discussions and comments. P.V. and M.A.M. acknowledge financial support from the Spanish "Ministerio de Ciencia, Innovación y Universidades" and the "Agencia Estatal de Investigación (AEI)" under Project Ref. PID2020-113681GB-I00 funded by MICIN/AEI/10.13039/501100011033.